toy giraffe was thrown horizontally from the edge of the table at a velocity of 7.5 m/s. if the table is 1.25 m high. what is the toys velocity just before it hits the ground
1
Expert's answer
2020-12-14T07:28:49-0500

Let's first find the time that the toy giraffe takes to reach the ground:


"y=\dfrac{1}{2}gt^2,""t=\sqrt{\dfrac{2y}{g}}=\sqrt{\dfrac{2\cdot 1.25\ m}{9.8\ \dfrac{m}{s^2}}}=0.505\ s."

Then, we can find the "y"-component of the toy's velocity (the "x"-component of the toy's velocity remains unchainged during the flight of the giraffe):


"v_y=-gt=-9.8\ \dfrac{m}{s^2}\cdot 0.505\ s=-4.95\ \dfrac{m}{s}."

The sign minus means that the vertical component of the toy's velocity directed downward.

Finally, we can calculate the toy's velocity just before it hits the ground:


"v=\sqrt{v_x^2+v_y^2}=\sqrt{(7.5\ \dfrac{m}{s})^2+(-4.95\ \dfrac{m}{s})^2}=8.98\ \dfrac{m}{s}."

Answer:

"v=8.98\ \dfrac{m}{s}."